A cylindrical wheel of uniform density with mass M=11kg and radius R= 2m rolls without slipping on an incline that makes an aMake sure to consider all parts of the problem with detail and accuracy. Please show your work and explain. Thank you!
Answer #1

solution F. B.D of block Ni Let block is sliding up with accoleration a. 11 30 30 using Necutons and law of 11 g Sinzo 11Now from eqh 11 g Sin 30 (T & T - 5,50) = lla 6 11 g Sinzo - 2T = 11a - 5.5a 11g sin 30 - 2T = 5.5a by 2 multiplying eqr adFor wheel, using equation of motion a = bit - I at2 a = 1.96 m/s2 Here, af = m I Fot7 X 1.9642 t² = 2XL 2 = 1.020 1.96 1.96 tDear friend if you have any doubts leave a comment... Thankyou.